(|(x-3)/2|)>=5
how can i solve this inequality?

Answer by lwsshak3(11628)

(|(x-3)/2|)>=5
how can i solve this inequality
**
|(x-3)/2|≥5
Two conditions make this inequality true
(x-3)/2≥=5
x-3≥10
x≥13
..
or
..
-[(x-3)/2]≥=5
multiply by (-1) and reverse inequality sign
(x-3)/2≤-5
x-3≤-10
x≤-7
number line:
<===]............[====>
.......-7.......... 13
solution:
(-∞,-7] U [13, ∞)